OpenStudy (anonymous):

Solve the problem as directed. The volume of a right circular cone varies jointly as the altitude and the square of the radius of the base. If the volume of the cone is 154 cu. in. when its altitude is 12 in. and the radius of the base is 3 1/2 in., find the altitude when the volume of the cone is 77 cu. in. and the radius of the base is 2 and 1/3 in. Altitude = _____a0 inches

OpenStudy (wolf1728):

Volume of a cone = (PI • r² •h) / 3 3 • volume / (PI • r²) = height or altitude height = 3 • 77 / (PI • 2.33333333²) height = 231 / (3.14159 • 5.4444444444) height = 13.5054337424

OpenStudy (wolf1728):

Well that would round to 13.51 Or to one decimal place 13.5